# Kramers nodal lines in intercalated TaS2 superconductors

### v1

@@ Abstract
Kramers degeneracy is one fundamental embodiment of the quantum mechanical nature of particles with half-integer spin under time reversal symmetry. Under the chiral and noncentrosymmetric achiral crystalline symmetries, Kramers degeneracy emerges respectively as topological quasiparticles of Weyl fermions and Kramers nodal lines, anchoring the Berry phase-related physics of electrons. However, an experimental demonstration for ideal Kramers nodal lines (KNLs) well isolated at the Fermi level is lacking. Here, we establish a class of noncentrosymmetric achiral intercalated transition metal dichalcogenide (TMD) superconductors with large Ising-type spin-orbit coupling, represented by InxTaS2, to host an ideal KNL phase. We provide evidence from angle-resolved photoemission spectroscopy (ARPES) with spin resolution, angle-dependent quantum oscillation (QO) measurements, and ab-initio calculations. Our work not only provides a realistic platform for realizing and tuning KNLs in layered materials, but also paves the way for exploring the interplay between KNLs and superconductivity, as well as applications pertaining to spintronics, valleytronics, and nonlinear transport.
@@
Department of Physics and Astronomy, Rice Center for Quantum Materials (RCQM), Rice University, Houston, Texas 77005, USA
@@ Introduction
Here, through comprehensive physical and thermodynamic properties characterization, angle-resolved photoemission spectroscopy (ARPES), and spin-resolved ARPES measurements, ab-initio calculations, and quantum oscillation measurements, we introduce a material class that exhibits an ideal KNL metallic phase, in the form of a noncentrosymmetric achiral intercalated transition metal dichalcogenide (TMD) family in the space group of $P\bar{6}m2$ [32, 33]. These are exemplified by InxTaS2 (x=1/2, 2/3, 1) (Fig. 1), all of which also feature superconducting ground states [34, 35, 36]. Due to the environment of the TMD layers (Figs. 1(a-b)), the KNL band topology is also coupled with spin-valley degrees of freedom (Figs. 1(c-d)). Such spin texture of KNLs in the intercalated TMD system under the $D_{3h}$ point group can be clearly visualized by the low energy band structure only considering spin-orbit coupling (SOC) terms (Fig. 1(d)), of which the momentum-dependent relativistic pseudospin splitting pattern is reminiscent of the recently generalized concept of nonrelativistic altermagnetism [37, 38]. However, the fact that the directional relativistic splitting is locked with the crystal structure symmetry, unlike the nonrelativistic altermagnetic splitting associated with antiferromagnetism, enables experimental investigation without complications of domain alignment. Furthermore, our theoretical analysis on the generic ideal KNL model at the Fermi level reveals several potential magnetic field-induced effects, including anomalous Hall effect (AHE) and Majorana zero modes that are theoretically possible in the $s$-wave superconducting phase. Additionally, the noncentrosymmetric achiral $P\bar{6}m2$ InxTaS2 associates the symmetries of the KNL material family with recent experimental reports of giant anomalous nonlinear transport [39] and strain-induced superconducting diode effect [40] in the isostructural PbTaSe2 with magnetic field-odd dependent behaviour.
@@ Results
The first step in constructing a noncentrosymmetric achiral crystal structure is to break the inversion symmetry. 2H-TaS2 and other TMDs of the 2H phase have been long known to have a centrosymmetric structure [41]. However, due to the local broken inversion symmetry constrained within a single layer, layer-resolved “hidden” spin polarization has been reported by surface sensitive ARPES techniques [42, 43]. In addition, extensive efforts focused on the monolayer and few-layer regime have shown evidence for unconventional Ising superconductivity that allows both spin-singlet and spin-triplet pairings [44, 45, 46, 47, 48, 49, 50, 51, 52]. To achieve inversion symmetry breaking in the bulk, we have intercalated 2H-TaS2 with trigonal indium layers, as shown in Figs. 1(a-b). The indium intercalation changes the 2H stacking to 1H stacking, and consequently, leads to broken inversion symmetry in the bulk of the InxTaS2 crystal structure. We also note that, as illustrated in Fig. 1(b), there is a structural difference between $x$ = 1/2, 2/3 and $x$ = 1: the In atoms are located above the Ta atoms for $x$ = 1/2 and 2/3 (left, Fig. 1b), while for $x$ = 1, the In atoms are above the S atoms (right, Fig. 1b). Such structural difference is supported by Rietveld refinements on the powder crystal XRD (see Supplementary Note 1). As a result, the Weyl nodal rings predicted in InTaS2 near $H$ point [35] (see Fig. 1(c) for the Brillouin zone (BZ) definition) based on a structure of In on top of Ta would instead become predicted Weyl points near $H$ [36]. Nonetheless, the KNL from the noncentrosymmetric achiral little group symmetry in InxTaS2 is fundamentally different from all the topological crossings predicted previously for similar compounds [35, 36, 53, 54]. Under the noncentrosymmetric achiral point group symmetry of InxTaS2, the electronic band structure from the two-band low energy Hamiltonian at $k_{z}$ = $\pi$ [6] only considering SOC terms shows an alternating pattern of pseudospin splitting, as shown in Fig. 1(d), while the black lines along $\Gamma-M$ indicate spin degeneracy and highlights the presence of the KNLs. The complete distribution of the KNLs is indicated by the purple lines in the first Brillouin zone (BZ) in Fig. 1(c). Furthermore, we visualize the realistic 3D Fermi surface (FS) of InTaS2 calculated from first-principles in the left panel of Fig. 1(e) showing the Kramers degeneracy along $\Gamma-M$ whenever the two FS sheets intersect. This can also be seen in the FS plotted for the $k_{z}$ = 0 plane where all Fermi pockets have pinched Kramers degeneracy when they cross the $\Gamma-M$ line (right of Fig. 1(e)).
@@
To demonstrate the physics of KNL and its spin-orbital texture, a candidate material of ideal KNL metals is highly desired. Here, “ideal KNL” is defined as an isolated KNL band that crosses $E_{\rm F}$, with a large SOC that results in a band splitting away from the KNL momentum directions. Among the noncentrosymmetric achiral intercalated TMD material family, we identify In1/2TaS2 to exhibit such ideal properties. First, to demonstrate the KNL properties, we carried out ab-initio calculations on the bulk Bloch spectral functions of In1/2TaS2, where the coherent potential approximation was used on the In sites accounting for the random vacancy without extending to a CDW supercell. As outlined by the red box in Fig. 2(a), the calculation shows a single set of isolated bulk KNL bands along $M-\Gamma-A-L$ crossing the Fermi level. Away from this momentum path, the KNL band splits into two spin-orbital branches, as shown along $\Gamma-K-M$ and $A-H-L$, where the $k_{z}$ = $\pi$ dispersions show the most prominent splitting. Projecting the Green’s function with the $\sigma_{z}$ operator in Fig. 2(b), it is unambiguously shown that the spin is strictly degenerate along the KNLs and highly polarized in $s_{z}$ along the off-KNL directions due to the Ising SOC. The splitting is most clearly demonstrated along $A-H-L$, but hybridized with In orbitals along $K-M$. To directly observe the ideal KNL band topology and provide the smoking-gun evidence for the spin splitting, we performed spin-integrated ARPES (Fig. 2) and spin-resolved ARPES (Fig. 3) measurements on In1/2TaS2. As shown in Fig. 2(c), the In1/2TaS2 FSs consist of two concentric pockets around $\bar{K}$ and $\bar{K\textquoteright}$ and two hexagonal pockets centered around $\bar{\Gamma}$, with no clear evidence of CDW folding or gaps. The ARPES band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ presented in Fig. 2(f) directly demonstrate the doubly-degenerate KNL crossing at $E_{\rm F}$ along $\bar{\Gamma}-\bar{M}$ and its clear splitting along $\bar{M}-\bar{K}-\bar{\Gamma}$. To theoretically capture the observed ARPES spectra with high accuracy, state-of-the-art one-step photoemission calculations for the FS and band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ are carried out (Figs. 2(d) and (g)), showing an excellent agreement with ARPES data, confirming the validity of the random vacancy modelling on the In sites. Furthermore, the spin-projected one-step calculation indicates a spin-polarized FS encoding spin-valley degree of freedom at $\bar{K}$ and $\bar{K\textquoteright}$ (Fig. 2(e)).
@@
The investigation of the spin-valley locking behavior associated with the KNLs using spin-resolved ARPES is presented in Fig. 3. As the spin polarization of the split bands should reverse on the two opposite directions away from the KNL, we can directly measure the spin polarization across a KNL. We carried out spin-resolved measurement of two pairs of momentum points on opposite sides of the $\Gamma-M$ KNL. $k_{1}$ and $k_{2}$ are along $\bar{K}-\bar{M}-\bar{K^{\prime}}$ (Fig. 3(a)). The spin up and down energy distribution curves (EDCs) selectively show the spin texture along the out-of-plane direction ($s_{z}$), and we plot them in direct comparison to the EDCs measured in spin-integrated mode in Fig. 3(b). Even though for both $k_{1}$ and $k_{2}$, the band closer to $E_{\rm F}$ exhibits weaker intensity as shown in the spin-integrated EDCs, the spin-resolved EDCs still show clear distinction for the split bands, namely for $k_{1}$ the lower binding energy band is dominantly spin up while the higher binding energy band is dominantly spin down. Furthermore, we can plot the $s_{z}$ polarization defined as $s_{z}=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}}$ (Fig. 3(c), $S$ being the Sherman function, $I_{\uparrow}$ and $I_{\downarrow}$ the spectral intensity measured in the spin up and spin down channels, respectively), which clearly shows the opposite spin polarization for the two bands split from the KNL. $k_{2}$, on the opposite side of the $\Gamma-M$ KNL, shows reversed spin polarization of the two split bands (Figs. 3(d) - (e)). To demonstrate this spin polarization reversal even more clearly, we show similar measurements for a pair of momentum points slightly away from the $\bar{K}-\bar{M}-\bar{K^{\prime}}$ high symmetry direction where the photoemission matrix elements allow comparable intensity of the two split bands as observed on each spin-integrated EDC ($k_{3}$ and $k_{4}$) (Figs. 3(f)-(j)). The spin-resolved ARPES data ($s_{z}$) measured here shows an even better defined peak structure and showcases consistent spin texture with spin polarization reaching as high as nearly 80%. Therefore, the combined experimental and theoretical results definitively demonstrate the coupled spin-valley polarization in In1/2TaS2, associated with the KNL band topology and the underlying Ising-type SOC shown previously only for monolayer TMD materials, which could potentially be relevant to applications in spintronics and valleytronics.
@@
As the noncentrosymmetric achiral symmetry is common across the family of intercalated TaS2, the KNL should be a universal property of all members of this family. In this section we present the evidence for KNLs for the stoichiometric InTaS2 in the $P\bar{6}m2$ space group. In addition, we also present the ARPES experimental measurement of the KNL band structure of the isostructural PbTaSe2 in the Supplementary Note 3. With full In intercalation, InTaS2 is more three-dimensional and exhibits stronger $k_{z}$ band dispersions when compared to In1/2TaS2 (Fig. 1(e)). Due to the $k_{z}$ broadening effect of the photoemission process, the measured FS of InTaS2 is more complex than the one of In1/2TaS2 (Fig. 4(a)). To account for potential surface states and the strong $k_{z}$ broadening effect in the vacuum ultraviolet ARPES spectra, we undertook ab-initio calculations projected to both In and S surface terminations. We find that the measured electronic structure can be best described by the sum of the calculations for both terminations. This is likely due to the fact that the In and S terminations have equal probabilities during the cleaving process. Therefore, the ARPES beam spot would probe the superposition of band structures under both terminations. The direct comparison of the measured and calculated FSs is shown in Figs. 4(a-b), where features such as the trefoil-like pattern centered at $\bar{K}$, the elliptical pockets centered at $\bar{M}$, and the $k_{z}$ broadened hexagonal pockets centered at $\bar{\Gamma}$ can be well reproduced. Importantly, we note that pairs of FSs are enforced to be degenerate whenever intersecting the $\Gamma-M$ direction, reflecting the properties of the KNL band topology. This can be further confirmed by the measured band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ (Fig. 4(c)), where we also overlay the bulk band calculations for $k_{z}$ = 0 and $\pi$. Clearly, both the bulk calculations and the ARPES data demonstrate that the bands along $\Gamma-M$ are doubly degenerate KNLs and split into two spin-orbital branches along $M-K-\Gamma$. To better capture the band dispersions observed in Fig. 4(c) beyond bulk contributions, we also display the surface calculations projected onto the In surface, which show an excellent match with the ARPES data.
@@
Since the FS of InTaS2 contains more 3D dispersions, as delineated in Fig. 1(e), and ARPES has significant spectral intensity contributions from $k_{z}$ broadening and surfaces, we use quantum oscillation (QO) measurements to provide a more complete investigation on the 3D bulk FS topology. The resistivity is measured as a function of the external magnetic field at different temperatures and field orientations, revealing Shubnikov de-Haas (SdH) oscillations (Figs. 4(d-i)). Strong SdH oscillations can be observed up to 20 K and 14 T after a smooth background subtraction, as can be seen in Fig. 4(d). The subsequent fast Fourier transform (FFT) on the SdH oscillations reveals that they consist of two frequencies Fα = 104 T and Fβ = 212 T (Fig. 4(e)). Despite the fact that Fβ is close to 2 $\times$ Fα, the effective mass fitting (inset of Fig. 4(e)) shows that the two frequencies originate from distinct Fermi surface cross-sectional areas, since the effective mass mβ is not twice that of mα. The small effective masses also imply that the underlying band dispersion is linear or close to linear.
@@
The complex Fermi surface of InTaS2, as illustrated in Fig. 1(e) and Figs. 4(a-c), should lead to QOs of various frequencies, while the SdH oscillations up to 14 T in Figs. 4(d-g) only consists of two frequencies. To further incorporate larger Fermi pockets, we measured the SdH oscillations up to 44.8 T at different magnetic field orientations (Figs. 4(h-i), supplemental information Figs. S5, 6). The SdH oscillations (supplemental information Fig. S5(a,b)) show a much more complex spectrum with more frequencies, consistent with the complex Fermi surface of InTaS2. Most interestingly, we discovered QOs of frequencies 6 kT and 12 kT (Fig. 4(h)). By comparing the experimental results (symbols) to DFT calculations (dashed lines) in Fig. 4(i), we found that these frequencies are related to the oscillations from the Fermi pocket $\epsilon$ (see the right panel of Fig. 1(e)) and its second harmonic. Such $\epsilon$ pockets encode pinched points enforced by both the $\Gamma-M$ and $\Gamma-A$ KNLs. Moreover, the frequency of the Fermi pocket $\epsilon$ (symbols) is significantly lower than the expected value for a cylindrical Fermi pocket (solid curve in Fig. 4(i)), which indicates that it must be a closed pocket as in the DFT calculation. We notice that at $\theta$ $\sim$ 9-10, the fundamental oscillations from the Fermi pocket $\epsilon$ vanishes while the second harmonic oscillations persist. This implies that at this field orientation, the quantum oscillations from the Fermi pocket $\epsilon$ are close to the spin-zero effect [56]. Overall, through ARPES and quantum oscillation measurements, we directly observe the Fermi pockets related to the KNLs in InTaS2. The underlying quasiparticles could contribute to the properties of InTaS2, for instance, the superconductivity.
Discussions and Conclusions

### v2

@@ Abstract
Kramers degeneracy is one fundamental embodiment of the quantum mechanical nature of particles with half-integer spin under time reversal symmetry. Under the chiral and noncentrosymmetric achiral crystalline symmetries, Kramers degeneracy emerges respectively as topological quasiparticles of Weyl fermions and Kramers nodal lines (KNLs), anchoring the Berry phase-related physics of electrons. However, an experimental demonstration for ideal KNLs well isolated at the Fermi level is lacking. Here, we establish a class of noncentrosymmetric achiral intercalated transition metal dichalcogenide superconductors with large Ising-type spin-orbit coupling, represented by InxTaS2, to host an ideal KNL phase. We provide evidence from angle-resolved photoemission spectroscopy with spin resolution, angle-dependent quantum oscillation measurements, and ab-initio calculations. Our work not only provides a realistic platform for realizing and tuning KNLs in layered materials, but also paves the way for exploring the interplay between KNLs and superconductivity, as well as applications pertaining to spintronics, valleytronics, and nonlinear transport.
@@
Department of Physics and Astronomy, Rice University, Houston, Texas 77005, USA
@@
Applied Physics Graduate Program, Smalley-Curl Institute, Rice University, Houston, Texas 77005, USA
@@
Physics Department, Florida State University, Tallahassee, FL 32306, USA
@@
Rice Center for Quantum Materials, Rice University, Houston, Texas 77005, USA
Department of Electrical and Computer Engineering, Rice University, Houston, Texas 77005, USA
Department of Materials Science and NanoEngineering, Rice University, Houston, Texas 77005, USA
Smalley-Curl Institute, Rice University, Houston, Texas 77005, USA
@@ Introduction
Here, through comprehensive physical and thermodynamic properties characterization, angle-resolved photoemission spectroscopy (ARPES), and spin-resolved ARPES measurements, ab-initio calculations, and quantum oscillation measurements, we introduce a material class that exhibits an ideal KNL metallic phase, in the form of a noncentrosymmetric achiral intercalated transition metal dichalcogenide (TMD) family in the space group of $P\bar{6}m2$ [32, 33]. These are exemplified by InxTaS2 ($x$=1/2 and 1) (Fig. 1), both of which also feature superconducting ground states [34, 35, 36]. In addition, similar physics of the KNLs in the isostructural intercalated TMD superconductors, In2/3TaS2 and PbTaSe2, is discussed in the Supplementary Information (notes 2 and 6). Due to the environment of the inversion symmetry broken TMD layers [37] (Figs. 1(a-b)), the KNL band topology is also coupled with spin-valley degrees of freedom (Figs. 1(c-d)). Such spin texture of KNLs in the intercalated TMD system under the D3h point group can be clearly visualized by the low-energy band structure only considering spin-orbit coupling (SOC) terms (Fig. 1(d)), of which the momentum-dependent relativistic pseudospin splitting pattern is reminiscent of the recently generalized concept of nonrelativistic altermagnetism [38, 39] but achieved without time reversal symmetry breaking. Meanwhile, the fact that the directional relativistic splitting is locked with the crystal structure symmetry, unlike the nonrelativistic altermagnetic splitting associated with the ligand-environment-enriched antiferromagnetism, enables experimental investigation without complications of domain alignment. Furthermore, our theoretical analysis on the generic ideal KNL model at the Fermi level reveals several potential magnetic field-induced effects, including anomalous Hall effect (AHE) and chiral Majorana modes that are theoretically possible in the $s$-wave superconducting phase. Additionally, the noncentrosymmetric achiral $P\bar{6}m2$ InxTaS2 associates the symmetries of the KNL material family with recent experimental reports of giant anomalous nonlinear transport [40] and strain-induced superconducting diode effect [41] in the isostructural PbTaSe2.
@@ Results
The first step in constructing a noncentrosymmetric achiral crystal structure is to break the inversion symmetry. 2H-TaS2 and other TMDs of the 2H phase have been long known to have a centrosymmetric structure [42] (see Fig. 1(a) for one exemplary inversion center). However, due to the local broken inversion symmetry constrained within a single layer, layer-resolved hidden spin polarization has been reported by surface-sensitive ARPES techniques [43, 44]. In addition, extensive efforts focused on the monolayer and few-layer regime have shown evidence for unconventional Ising superconductivity that allows both spin-singlet and spin-triplet pairings [45, 46, 47, 48, 49, 50, 51]. To achieve inversion symmetry breaking in the bulk, we have intercalated 2H-TaS2 with trigonal indium layers, as shown in Figs. 1(a-b). The indium intercalation changes the 2H stacking to 1H stacking, and consequently, leads to broken inversion symmetry in the bulk of the InxTaS2 crystal structure. We also note that, as illustrated in Fig. 1(b), there is a structural difference between $x$ = 1/2, 2/3 and $x$ = 1: the In atoms are located above the Ta atoms for $x$ = 1/2 and 2/3 (left, Fig. 1b), while for $x$ = 1, the In atoms are above the S atoms (right, Fig. 1b). Such structural difference is supported by Rietveld refinements on the powder crystal XRD (see Supplementary Note 1). As a result, the Weyl nodal rings predicted in InTaS2 near $H$ point [35] (see Fig. 1(c) for the Brillouin zone (BZ) definition) based on a structure of In on top of Ta would instead become predicted Weyl points near $H$ [36]. Nonetheless, the KNL from the noncentrosymmetric achiral little group symmetry in InxTaS2 is fundamentally different from all the topological crossings predicted previously for similar compounds [35, 36, 52, 53]. In Fig. 1(b), we highlight the most important symmetry operation for the formation of the KNLs in the InxTaS2 family as magenta planes, the $M_{z}$ mirror. Under the point group symmetry of InxTaS2, the electronic band structure from a minimal two-band low-energy Hamiltonian at $k_{z}=0$ [6] only considering SOC terms shows an alternating pattern of pseudospin splitting, as shown in Fig. 1(d), while the green tubes represent the topologically nontrivial KNLs. The complete distribution of the KNLs is indicated by the purple lines in the first Brillouin zone (BZ) in Fig. 1(c). Furthermore, we visualize the realistic Fermi surface (FS) Bloch spectral function (BSF) of In1/2TaS2 calculated from first-principles at $k_{z}=0$ and $k_{z}=\pi$ in Fig. 1(e) (right) showing the Kramers degeneracy along $\Gamma-M$ and $A-L$ whenever the two hexagonal-like FS sheets centered around $k_{x}=k_{y}=0$ intersect. The three-dimensional (3D) FSs extracted from high intensities of the BSF cast in voxel-style plot in Fig. 1(e) (left) reflect the quasi-two-dimensional (2D) nature of In1/2TaS2 at the Fermi level.
@@
To demonstrate the physics of KNL and its spin-orbital texture, a candidate material of ideal KNL metals is highly desired. Here, “ideal KNL” is defined as an isolated KNL band that crosses $E_{\rm F}$, with a large SOC that results in a band splitting away from the KNL momentum directions. Among the noncentrosymmetric achiral intercalated TMD material family, we identify In1/2TaS2 to exhibit such ideal properties. First, to demonstrate the KNL properties, we carried out ab-initio calculations on the bulk BSF of In1/2TaS2, where the coherent potential approximation was used on the In sites accounting for the random vacancy without extending to a CDW supercell. As outlined by the red box in Fig. 2(a), the calculation shows a single set of isolated bulk KNL bands along $M-\Gamma-A-L$ crossing the Fermi level. Away from this momentum path, the KNL band splits into two spin-orbital branches, as shown along $\Gamma-K-M$ and $A-H-L$, where the $k_{z}$ = $\pi$ dispersions show the most prominent splitting. Projecting the Green’s function with the $\sigma_{z}$ operator in Fig. 2(b), it is unambiguously shown that the spin is strictly degenerate along the KNLs and highly polarized in $s_{z}$ along the off-KNL directions due to the Ising SOC. The splitting is most clearly demonstrated along $A-H-L$, but hybridized with In orbitals along $K-M$. To directly observe the ideal KNL band topology and provide the smoking-gun evidence for the spin splitting, we performed spin-integrated ARPES (Fig. 2) and spin-resolved ARPES (Fig. 3) measurements on In1/2TaS2, of which the same sample was exfoliated and examined by scanning electron microscopy with energy dispersive x-ray spectroscopy to confirm spatial homogeneity and the In concentration (see Supplementary Note 3). As shown in Fig. 2(c), the In1/2TaS2 FSs consist of two concentric pockets around $\bar{K}$ and $\bar{K\textquoteright}$ and two hexagonal pockets centered around $\bar{\Gamma}$, with no clear evidence of CDW folding or gaps. The ARPES band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ presented in Fig. 2(f) directly demonstrate the doubly-degenerate KNL crossing at $E_{\rm F}$ along $\bar{\Gamma}-\bar{M}$ and its clear splitting along $\bar{M}-\bar{K}-\bar{\Gamma}$. The chosen 75 eV of photons here is justified to be close to the $k_{z}=0$ plane through a photon-energy-dependent scan focusing on certain $k_{z}$ sensitive features at deeper binding energies, as presented in Supplementary Fig. 4 of the Supplementary Note 4. To theoretically capture the observed ARPES spectra with high accuracy, state-of-the-art one-step photoemission calculations for the FS and band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ are carried out with the In termination (see Figs. 2(d) and (g)), showing an excellent agreement with ARPES data, confirming the validity of the random vacancy modelling on the In sites. An in-depth discussion for the termination dependent band structure of In1/2TaS2 is presented in Supplementary Fig. 7 of the Supplementary Note 5, supporting the adopted In termination. Furthermore, the spin-projected one-step calculation indicates a spin-polarized FS encoding spin-valley degree of freedom at $\bar{K}$ and $\bar{K\textquoteright}$ (Fig. 2(e)).
@@
In the In1/2TaS2 family, the spin-valley polarization arises from the inversion symmetry breaking [37] which is satisfied as a prerequisite of the KNL little group symmetries. Meanwhile, the mirror symmetries such as the magenta ones denoted in Fig. 1(b) in the noncentrosymmetric achiral little group generate KNLs concentrating Berry curvature and forcing spin degeneracy robust against SOC. Further, heavy elements such as Ta provide strong strength of the SOC, exhibiting an Ising-type splitting at valleys such as $K$ and $K^{\prime}$ mainly due to the broken $\Gamma-K-H-A$ mirror [54]. Experimentally, the investigation of the spin-valley locking behavior associated with the KNLs using spin-resolved ARPES is presented in Fig. 3. As the spin polarization of the split bands should reverse on the two opposite directions away from the KNL, we can directly measure the spin polarization across a KNL. We carried out spin-resolved measurement of two pairs of momentum points on opposite sides of the $\Gamma-M$ KNL. $k_{\mathrm{1}}$ and $k_{\mathrm{2}}$ are along $\bar{K}-\bar{M}-\bar{K^{\prime}}$ (Fig. 3(a)). The spin up and down energy distribution curves (EDCs) selectively show the spin texture along the out-of-plane direction ($s_{z}$), and we plot them in direct comparison to the EDCs measured in spin-integrated mode in Fig. 3(b). Even though for both $k_{\mathrm{1}}$ and $k_{\mathrm{2}}$, the band closer to $E_{\rm F}$ exhibits weaker intensity as shown in the spin-integrated EDCs, the spin-resolved EDCs still show clear distinction for the split bands, namely for $k_{\mathrm{1}}$ the lower binding energy band is dominantly spin up while the higher binding energy band is dominantly spin down. Furthermore, we can plot the $s_{z}$ polarization defined as $P_{z}=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}}$ (Fig. 3(c), $S$ being the Sherman function, $I_{\uparrow}$ and $I_{\downarrow}$ the spectral intensity measured in the spin up and spin down channels, respectively), which clearly shows the opposite spin polarization for the two bands split from the KNL. $k_{\mathrm{2}}$, on the opposite side of the $\Gamma-M$ KNL, shows reversed spin polarization of the two split bands (Figs. 3(d) - (e)). To demonstrate this spin polarization reversal even more clearly, we show similar measurements for a pair of momentum points slightly away from the $\bar{K}-\bar{M}-\bar{K^{\prime}}$ high symmetry direction where the photoemission matrix elements allow comparable intensity of the two split bands as observed on each spin-integrated EDC ($k_{\mathrm{3}}$ and $k_{\mathrm{4}}$) (Figs. 3(f)-(j)). The spin-resolved ARPES data ($s_{z}$) measured here shows an even better defined peak structure and showcases consistent spin texture with spin polarization reaching as high as nearly 80%. Therefore, the combined experimental and theoretical results definitively demonstrate the coupled spin-valley polarization in In1/2TaS2, associated with the KNL band topology and the underlying Ising-type SOC shown previously only for monolayer TMD materials, which could potentially be relevant to applications in spintronics and valleytronics.
@@
As the noncentrosymmetric achiral symmetry is common across the family of intercalated TaS2, the KNL should be a universal property of all members of this family. In this section we present the evidence for KNLs for the stoichiometric InTaS2 in the $P\bar{6}m2$ space group. In addition, we also present the ARPES experimental measurement of the KNL band structure of the isostructural PbTaSe2 in the Supplementary Note 6. With full In intercalation, InTaS2 is more three-dimensional and exhibits stronger $k_{z}$ band dispersions when compared to In1/2TaS2 (Fig. 1(e)). Due to the $k_{z}$ broadening effect of the photoemission process, the measured FS of InTaS2 (Fig. 4(a)) is more complex than the one of In1/2TaS2. Meanwhile, stronger $k_{z}$ integration in the photoemission process of InTaS2 gives out weaker photon-energy-dependence for the band dispersions, as evidenced by the $h\nu$-scan along $\bar{\Gamma}-\bar{M}$ shown in Supplementary Fig. 5. To account for potential surface states and the strong $k_{z}$ broadening effect in the vacuum ultraviolet ARPES spectra, we undertook ab-initio calculations projected to both In and S surface terminations. We find that the measured electronic structure in Fig. 4(a) can be best described by the summation of the calculations for both terminations. This is likely due to the fact that the In and S terminations have equal probabilities during the cleaving process. Therefore, the ARPES beam spot would probe the superposition of band structures under both terminations, while the separate contributions from each termination are elaborated in Supplementary Fig. 6 of Supplementary Note 5. The direct comparison of the measured and calculated FSs is shown in Figs. 4(a-b), where features such as the trefoil-like pattern centered at $\bar{K}$, the elliptical pockets centered at $\bar{M}$, and the $k_{z}$ broadened hexagonal pockets centered at $\bar{\Gamma}$ can be well reproduced. Importantly, we note that pairs of FSs are enforced to be degenerate whenever intersecting the $\Gamma-M$ direction, reflecting the properties of the KNL band topology. This can be further confirmed by the measured band dispersions along $\bar{\Gamma}-\bar{M}-\bar{K}-\bar{\Gamma}$ (Fig. 4(c)), where we also overlay the bulk band calculations for $k_{z}$ = 0 and $\pi$. Clearly, both the bulk calculations and the ARPES data demonstrate that the bands along $\Gamma-M$ are doubly degenerate KNLs and split into two spin-orbital branches along $M-K-\Gamma$. To better capture the band dispersions observed in Fig. 4(c) beyond bulk contributions, we also display the surface calculations projected onto the In surface in Fig. 4(d), which show an excellent match with the ARPES data.
@@
Since the FS of InTaS2 contains more 3D dispersions, as delineated in Fig. 5(a), and ARPES has significant spectral intensity contributions from $k_{z}$ broadening and surfaces, we use quantum oscillation (QO) measurements to provide a more complete investigation on the 3D bulk FS topology. The assignment of the Fermi pockets discussed in QO measurements and density functional theory (DFT) calculations is illustrated in Fig. 5(b) for a cross section of the FS at $k_{z}=0$. Next in Figs. 5(c-h), the resistivity is measured as a function of the external magnetic field at different temperatures and field orientations, revealing Shubnikov de-Haas (SdH) oscillations. Strong SdH oscillations can be observed up to 20 K and 14 T after a smooth background subtraction, as can be seen in Fig. 5(c). The subsequent fast Fourier transform (FFT) on the SdH oscillations reveals that they consist of two frequencies Fα = 104 T and Fβ = 212 T (Fig. 5(d)). Despite the fact that Fβ is close to 2 $\times$ Fα, the effective mass fitting (inset of Fig. 5(d)) shows that the two frequencies originate from distinct Fermi surface cross-sectional areas, since the effective mass mβ is not twice that of mα. The small effective masses also imply that the underlying band dispersion is linear or close to linear.
@@
The complex Fermi surface of InTaS2, as illustrated in Figs. 5(a-b) and Figs. 4(a-d), should lead to QOs of various frequencies, while the SdH oscillations up to 14 T in Figs. 5(c-f) only consist of two frequencies. To further incorporate larger Fermi pockets, we measured the SdH oscillations up to 44.8 T at different magnetic field orientations (Figs. 5(g-h), Supplementary Information Figs. 10, 11). The SdH oscillations (Supplementary Information Fig. 10(a,b)) show a much more complex spectrum with more frequencies, consistent with the complex Fermi surface of InTaS2. Most interestingly, we discovered QOs of frequencies 6 kT and 12 kT (Fig. 5(g)). By comparing the experimental results (symbols) to DFT calculations (dashed lines) in Fig. 5(h), we found that these frequencies are related to the oscillations from the Fermi pocket $\epsilon$ (see Fig. 5(b)) and its second harmonic. Such $\epsilon$ pockets encode pinched points enforced by both the $\Gamma-M$ and $\Gamma-A$ KNLs. Moreover, the frequency of the Fermi pocket $\epsilon$ (symbols) is significantly lower than the expected value for a cylindrical Fermi pocket (solid curve in Fig. 5(h)), which indicates that it must be a closed pocket as in the DFT calculation. We notice that at $\theta$ $\sim$ 9-10, the fundamental oscillations from the Fermi pocket $\epsilon$ vanish while the second harmonic oscillations persist. This implies that at this field orientation, the quantum oscillations from the Fermi pocket $\epsilon$ are close to the spin-zero effect [55]. Overall, through ARPES and quantum oscillation measurements, we directly observe the Fermi pockets related to the KNLs in InTaS2. The underlying quasiparticles could contribute to the properties of InTaS2, for instance, the superconductivity.
Discussion